OpenStudy (anonymous):

How do you find whether or not a coordinate (don't have a specific example) lies on a particular line? Also, what does Y = MX+ C stand for?

OpenStudy (anonymous):

y=mx+b is the equation of a straight line or y=mx+c, must be in the UK yeah? :P When you use y = mx+b to graph, the y represents how far up the x represents how far long m is the slope (steepness of the line) b is the y - intercept, where the line crosses the y - axis  This is an example of where m (slope) = 2 \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] Basically saying, \[m = \frac{ Change ~ In ~ Y }{ Change ~ In ~ X }\] and you can see in the graph, that b = 1 since that's where the line crosses the y - axis. In y = mx+ b form your solution would be y=2x+1 Another common example, is find the equation of a straight line that has a slope m = 4 and passes though the point (-1, -6). Once again we look at y = mx+b formula, and we see we need to find b (y - intercept). -6=4*-1+b -6=-4+b -2 = b Once you find b you can use y = mx+b form again and put it all together so you get y = 4x- 2 Another way to solve a problem like this would be to use the point - slope form. \[y-y _{1}=m(x-x _{1})\] Where the \[(x _{1}, y _{1})\] would represent your coordinate. so you'd get, y - (-6) = 4(x-(-1)) y + 6 = 4(x+1) y+6 = 4x+4 y = 4x+4-6 y = 4x-2 You can see that it's similar when we were using y = mx+b form. Well I hope that helps you somewhat, take care :)!